OpenStudy (anonymous):

Jonathan sells 34 of his used CDs for $8.49 each. He also sells 13 vinyl records for $3.15 each. He gets a 20% bonus for the cost of the records because they are unscratched. How much did Jonathan earn from selling his used CDs and records?

OpenStudy (anonymous):

Ok so first lets find the total of the CDs/ 8.49 * 34 Next we will calculate the vinyl. 3.15 *13 Let x represent the total of the cds and y represent the total of the vinyl after his bonous. After solving we would add x+y

OpenStudy (anonymous):

i came up with this equation : 34 * 8.49 + (13 * 3.15) - .2 ; 288.66 + 40.95

OpenStudy (anonymous):

So 8.49 = 288.66 3.15*13 = 40.95 40.95* .2 = 8.19 40.95 + 8.19 = 49.14 288.66+49.14 = 337.8 So he earns $337.80

OpenStudy (anonymous):

Ohh my gosh I totally forgot to multiply by .2 im sorry

OpenStudy (anonymous):

Haha well do you understand what I did?

OpenStudy (anonymous):

yes i do